A solution of sodium acetate (NaCH3COO)has a pH of 9.73. The acid-dissociation constant for acetic acid is 1.8×10−5 What is the molarity of the solution?

Express your answer to two significant figures and include the appropriate units.
Answer #1

pKa = - log Ka = - log (1.8 * 10^-5) = 4.74

NaCH3COO is the salt of weak acid and strong base.

thus

salt hydrolysis occurs.

so Using

pH = 1/2 * [pKw + pKa + log C]

or

9.73 = 1/2 * [14 + 4.74 + log C ]

or

C = 5.2 M

thus

molarity of the solution = 5.2 M